9) Use your value of x to calculate a value for 𝑒𝑥 from the first four

terms.

X = 1



1
Expert's answer
2022-04-18T00:31:47-0400

"e^x=\sum_{k=0}^{\infty}\frac{x^k}{k!}"

First 4 terms:

"e^x\approx 1+x+\frac{x^2}{2}+\frac{x^3}{6}"

For x = 1:

"e^1=e\approx1+1+\frac{1}{2}+\frac{1}{6}=\frac{16}{6}\approx2.67"